OpenStudy (anonymous):

newton's universal law of gravitation asserts that two point masses m1 and m2 seperated by a distance s, attract one another with a force F=(km1m2)/s^2 , k being a universal constant.

OpenStudy (anonymous):

wats the main objective of this question?

OpenStudy (anonymous):

sorry never got to finish if s=16 meter, use differentials to estimate the change in s that will decrease the force F by 10%

OpenStudy (anonymous):

So, find ds : dF = 0.1F?

OpenStudy (anonymous):

sorry super tired so use related rates? dF is 0.1F and differentiate and solve for ds?

OpenStudy (anonymous):

That's what I'm thinking, but I'm not sure if I'm setting it up right. I scribbled out some stuff and got a negative value for ds, and that can't be right because decreasing s will increase F...

OpenStudy (anonymous):

I think I need to put in dF = -0.1F . . . maybe. I'm a bit tired too. X-D

OpenStudy (anonymous):

ya a neg 0.1 makes more sense to me too

OpenStudy (anonymous):

I don't even think this is a related rates thing. It's more like using partials to find marginal cost or %error or something like that.

OpenStudy (anonymous):

ok good cuz i was stuck thinking how to take the derivative of km1m2..

OpenStudy (anonymous):

I'm assuming km1m2 stays constant.

OpenStudy (anonymous):

I'm getting what I think is a reasonable answer. I'm going to test it with some arbitrary values to see if it's in the ballpark.

OpenStudy (anonymous):

I let C=km1m2, so F=C*s^-2 then dF/ds = -2C*s^-3 etc.

OpenStudy (anonymous):

my first attempt was F-(F*0.1)=(km1m2)/s^2 so if i substitute C in I getcha ok thanks!

OpenStudy (anonymous):

Ok, this does seem like the right way to go. I tested it against the actual difference and the estimate is accurate to 0.8% Solve for ds and use dF=-0.1F.

OpenStudy (anonymous):

ok i feel pretty dumb in asking but am i completely off track thinking of -0.1F=-2C*s^-3 ?

OpenStudy (anonymous):

There does seem to be something missing...

OpenStudy (anonymous):

lol please don't say between my ears

OpenStudy (anonymous):

LOL. I can neither confirm nor deny that proposition. No, you're missing the ds at the end.

OpenStudy (anonymous):

haha ok thanks

OpenStudy (anonymous):

doing it that way do you get a final change of .8?

OpenStudy (anonymous):

wait going back to physics if the distance between the two masses... (s) increase the gravitational force decreases by the same factor right? so if you want the force to decrease by a factor of 0.10 then you would increase the distance by a factor of 0.10.. right?

OpenStudy (anonymous):